OWG Meeting 13-Mar-2013

From GlueXWiki
Jump to: navigation, search


  • Announcements
    • Labview FPGA boards as alternative to CAEN V1495 boards
    • Still need to arrange for Ed to talk more on FPGA programming
  • Review of minutes from 27-Feb-2013 meeting
  • Front-end module data format (uncompressed) - Dave A, Elliott
  • What online projects can other groups take over? - All
  • DAQ testing in Rack Room - Dave A, Elliott
  • Temporary switch in Hall for summer tests - Elliott
  • Rebaseline status - Hovanes, Dave L, Elliott
  • Counting House installation update - Hovanes, Elliott
  • Hall D Computer Cluster (HDCC) status - Hovanes, Elliott
  • Status of processing uncompressed and compressed simulated raw data and monitoring/L3 farm - Dave L, Elliott
  • Solenoid test update - Elliott
  • EPICS update - Hovanes
  • Controls update - Elliott, Hovanes
  • JInventory database - Elliott
  • DAQ group projects- Bryan, Dave A, Graham, Carl, Vardan, Elliott
    • EVIO, cMsg, CODA3, ECS/CSS, disentangler, readout list, Java/cMsg GUI
  • Electronics, trigger and DAQ status reports - Chris, Alex, Dave, Fernando, Ben, Ed, William, Bryan
  • Mantis Task Tracker
  • Midas Elog

  • Calibration/Alignment issues?


1:30 PM Wed 13-Mar-2013 CC F326


Next Meeting

1:30 PM Wed 27-Mar-2013 CC F326


Present: Elliott W, Simon T, Dave L, Bryan M, Hovanes E, Dave A, Yi Q, Beni Z.

Front-end Data Format

We how have agreement that all FPGA boards read out via VME will comply with Dave A and Ed J's note on EVIO data format. The latest revision addressed all of Ben's concerns. It may take a while for this to be implemented in some of the board, but it should be implemented by the time we start reading them out. We further asked Graham to place the note in some official place that we can refer to. Now we need a similar agreement concerning compressed EVIO format. Dave A notes the latter may not be possible at this stage.

Online Projects

Elliott presented a list of online projects that outside groups might potentially work on (see list above and on the Online Wiki page). Four categories were identified, high and low priority, and projects that can be done remotely vs locally. Some projects on the list may not be in the proper category. Elliott noted that he was pushing the envelope on what might be done remotely in hopes of luring people in. Although we went over the list quickly, there was much discussion on whether we need an event reorder stage in the DAQ event pipeline. No conclusions, but no compelling arguments either way.

DAQ Test

Dave A reported on the status of setting up in the Rack room. Brent installed the switch in the left side. Infiniband currently being installed, all cards are here. We will be swapping out the 4 x 16 core box for a number of other more suitable nodes from Chip next week. Dave is currently using one mini-VME crate, we need to get full crates in when Fernando is done with them. Dave is in the process of installing CODA3 on our file server. With luck installation should be done next week and testing should commence.

Temporary Switch in Hall

Elliott reported that we'll need a temporary switch in the hall by the time we start detector testing this summer until the main fiber trunk is installed in the fall. The switch will connect to the main switch in the Rack room through a conduit and will be mounted on the North wall at a location to be chosen by Tom. All production switches in the DAQ racks in the hall will connect to this switch until the fibers are installed, at which time it will be removed. Plenum-rated cable (expensive!) must be used to connect to the main switch (since it runs under a floor), CL2/3 cable can be used in the hall itself.


Hovanes reported that our work is nearly done and that the exercise officially ends on 26-Mar. Most equipment lines are unchanged, most labor lines are completely rewritten. Many start/stop dates were changed.

Counting House

  • PR for furniture held up in Facilities (we need this before summer).
  • Facilities will run power from Rack room wall to consoles.
  • Cable tray installed in Rack room.
  • Almost all computers installed in Rack room.
  • File server configured, accounts created, RAID working.
  • CUE disks mounted.

Solenoid Test

Elliott reported that the compressors/cold boxes have been repaired and are beginning to cool down. Most recent problem was a carbon dust blockage in a pipe due to disintegration of the carbon in a filter. Optimistic scenario has them making liquid in about a week, then a week of magnet cooldown and we can begin powered operations. Unfortunately the Danfysik power supply failed to ramp to full current during commissioning, current thinking is that power transistors got damaged while it was in Hall A. Unfortunately there are over 1300 power transistors and many are difficult to replace. Mark and co. are replacing some to see if it helps.


Beni and Yi have been working on some gui's and drivers for the HV/LV system. We agreed with Accelerator to purchase computers they need to implement their MyA archiver for Hall D.


Dan and John presented their latest thoughts on controls for the FCAL darkroom at the last meeting. They need suggestions on which sensors to purchase as they prefer to use ones in common use at JLab. Mark and Scot will advise them.


Alexandre Comsonne from Hall A is interested in adopting the system, and is willing to work on improving the GUI code. We are still waiting for Sergey B's cable database feature. Hovanes asked whether the system will be capable of generating translation tables, supplying information for trigger configurations, etc, and if so, will it be implemented in time. Hovanes and others noted that it is time to get our act together concerning conventions for using JInventory. Someone noted that there are too many fields to fill in for new entries, but Elliott explained that very few are mandatory, and that if a field doesn't make sense for a particular item then just leave it empty (e.g. the "Telecom" field for a building).